OpenStudy (samanthagreer):

1. Which of the following equations has exactly one real solution? A. 5(x + 10) = 5x + 50 B. 7x + 10 = -5x - 18 C. 7x + 10 = 7x - 18 D. 5x + 10 = 5x + 10

OpenStudy (anonymous):

What do you think it is? @samanthagreer

OpenStudy (anonymous):

Give me your best guess.

OpenStudy (samanthagreer):

d

OpenStudy (danjs):

the one that you can solve for 1 sol

OpenStudy (danjs):

x=number

OpenStudy (anonymous):

Okay well lets work on this @samanthagreer

OpenStudy (anonymous):

Lets first start with A.

OpenStudy (anonymous):

Lets try and solve it by distributing 5(x+10)

OpenStudy (anonymous):

Do you know how to distribute @samanthagreer

OpenStudy (danjs):

A is the same thing on both sides, you get a true statement, x can be anything, infinite solutions

OpenStudy (samanthagreer):

multiply 5 times what is in the perentaces

OpenStudy (anonymous):

Okay yes so when you do that 5(x+10) would become 5x+50! So now what is the answer?

OpenStudy (samanthagreer):

55x

OpenStudy (danjs):

B looks like you can solve for x= C turns into a False statement, or No solutions

OpenStudy (anonymous):

@samanthagreer so using the info I just gave you what is the answeR?

OpenStudy (anonymous):

@DanJS you got this lol

OpenStudy (samanthagreer):

A

OpenStudy (danjs):

each of these is sort of like solving a system of two linear equations y = mx+b, setting them equal will tell you what x values may have a point where the two lines cross and a solution`

OpenStudy (danjs):

like A) is a line y = 5(x+10) and another y = 5x+50 they are the same line, think of them as overlapping each other at an infinite number of points, solutions

OpenStudy (samanthagreer):

wait its b

OpenStudy (samanthagreer):

right

OpenStudy (danjs):

those two intersect , the solution is a value for x for that intersection point

OpenStudy (danjs):

yes, you can solve b) for a value of x

OpenStudy (samanthagreer):

ok thanks for the hel pi will give you a medal if you give tom one

OpenStudy (danjs):

C) gives you 10=-18, false, the two lines are parallel, never intersecting,

OpenStudy (danjs):

no solutions in that case

OpenStudy (danjs):

welcome